OpenStudy (agentjamesbond007):

Write y=e^(2x-x^2) in terms of x=

myininaya (myininaya):

What is the inverse of the exponential function?

OpenStudy (agentjamesbond007):

ln?

myininaya (myininaya):

Ok so do ln( ) of both sides and what do we get?

OpenStudy (agentjamesbond007):

ln y = 2x - x^2

myininaya (myininaya):

Here is another hint: You will have to use quadratic formula :)

myininaya (myininaya):

See if you can see why

sam (.sam.):

y=e^(2x-x^2) lny=2x-x^2 Complete the square lny=-x^2+2x lny=-(x-1)^2+1 Can you solve from here?

myininaya (myininaya):

Or what sam did that works :)

OpenStudy (agentjamesbond007):

Thank you